OpenStudy (anonymous):

how can i rewrite this 8x-2y=0 using the formula of f(x)=mx+b

OpenStudy (anonymous):

you can subtract 8x from both sides to start

OpenStudy (anonymous):

you need to get y by itself so you the divide both sides by -2

OpenStudy (anonymous):

then u simplify

OpenStudy (anonymous):

do u still need help bud?

OpenStudy (anonymous):

it is complicated for me xD cause i dont know how to do that xD i hate math xD

OpenStudy (anonymous):

can u give me the equation so i can get the idea to solve the other equation

OpenStudy (anonymous):

sure f(x)=-4x

OpenStudy (anonymous):

u see how i got that?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

8x-2y=0 8x-2y-8x=0-8x -2y=-8x 2y=8x y=4x

OpenStudy (anonymous):

yup thats it

OpenStudy (yanasidlinskiy):

By the way, Everyone, \(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)

OpenStudy (anonymous):

hey yo, dont forget to give the medal atleast (y)